If f(x)=1-3x and g(x)=1/x2+1,find f(g(0)) and g(f(1)).
(by 1/x2+1 i mean 1 over x squared plus 1) PLEASE HELP!!!








Answer by stanbon(75887)

If f(x)=1-3x and g(x)=1/(x^2+1),find f(g(0)) and g(f(1))
-------------------------------
f[g(o)] = f[1] = 1-3 = -2
-----------------------------------
g[f(1)] = g[-2] = 1/(2^2+1) = 1/5
=========================================
Cheers,
Stan H.